Mobiui ravvice quarunca DE x(y2 + up – y(x2 + ung = (x2 - y2)u

Find the general solution of the quasilinear PDE
Answer #1

O Quasilinear equation xt -+u) p —(x*4 ч) 1 - (-?) и the characteritic equation are doe dy , x (32+4) du (22) — 5-9 (+4) firsIndt udat ydy = du integrating both of sides A + y2 = ured ~ 22+42 = 24+Cg zu be+y2–24 = (270 The general solution is given b